OpenStudy (anonymous):

true/false: For the polynomial f(x)=1-2x-3x^4, as x approaches infinity, f(x) approaches negative infinity.

OpenStudy (anonymous):

Whenever you evaluate limits, our first urge is to directly substitute the value into the function. So we see that \(\displaystyle\lim\limits_{x\to\infty} 1-2x-3x^4 = 1-2\infty-3\infty^4 = 1-\infty-\infty\). What is that final expression equivalent to?

OpenStudy (anonymous):

- infinity?

OpenStudy (anonymous):

Correct. So is the statement true or false? :-)

OpenStudy (anonymous):

true ?

OpenStudy (anonymous):

Yes, correct. Infinity can be a tricky thing to understand...does this explanation make any sense at all?

OpenStudy (anonymous):

yeah, thank you :)